40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Jul 14 00:57:11 2009, Alberto, Update, Locking, photodiode alignment check 
    Reply  Wed Jul 15 12:14:08 2009, Alberto, Update, Locking, photodiode alignment check 2009-07-15_PD9spectrumPDF.pdf
       Reply  Thu Jul 16 01:00:56 2009, Alberto, Update, Locking, PD9 aligned 2009-07-15_PD9spectrumPDF02.pdf
Message ID: 1749     Entry time: Wed Jul 15 12:14:08 2009     In reply to: 1742     Reply to this: 1755
Author: Alberto 
Type: Update 
Category: Locking 
Subject: photodiode alignment check 

Quote:

Since lately the alignment of the input beam to the interferometer has changed, I went checking the alignment of the beam on the photodiodea. They were all fine except for pd9, that is AS DD 199. Here the DC is totally null. The beam seems to go right on the diode but the scope on the PD's DC output shows no power. This is really strange and bad.

After inspecting PD9 with the viewer and the cards, the beam looks like it is aligned to the photodiode althought there is no signal at the DC output of the photodetector. So I checked the spectrum for PD9_i and Q (see attachments) and it seems that those channels are actually seeing the beam. I'm going to check the alignemtn again and see the efefct on the spectra to make sure that the beam is really hitting the PD.

 

Attachment 1: 2009-07-15_PD9spectrumPDF.pdf  33 kB  | Hide | Hide all
2009-07-15_PD9spectrumPDF.pdf
ELOG V3.1.3-